### v1

# Behavioral Manifolds: Representing the Landscape of Grasp Affordances in the Relative Pose Space

Abstract
The use of machine learning to investigate grasp affordances has received extensive attention over the past several decades. The existing literature provides a robust basis to build upon, though a number of aspects may be improved. Results commonly work in terms of grasp configuration, with little consideration for the manner in which the grasp may be (re-)produced from a reachability and trajectory planning perspective. In addition, the majority of existing learning approaches focus of producing a single viable grasp, offering little transparency on how the result was reached, or insights on its robustness. We propose a different perspective on grasp affordance learning, explicitly accounting for grasp synthesis; that is, the manner in which manipulator kinematics are used to allow materialization of grasps. The approach allows to explicitly map the grasp policy space in terms of generated grasp types and associated grasp quality. Results of numerical simulations illustrate merit of the method and highlight the manner in which it may promote a greater degree of explainability for otherwise intransparent reinforcement processes.
Figure 1: Considered end-effector (Robotiq 3-digit gripper), shown example behavioral manifolds for a hand-drill (from the YCB dataset) for the considered gripper and grasp synthesis policy discussed in Section III (left); grasping a cylinder, actuated joints, centers of mass of the cylinder and gripper base, and contact points shown in blue (right).
I Introduction
The use of machine learning in robotic automation is dramatically impacting artificial agents’ capabilities, accelerating learning and skill development capacities, providing perspectives of productive artificial cognitive systems in use on factory floors ([1]). A meaningful dimension in these developments is the growing use of Reinforcement Learning (RL) to support robotic manipulators in learning to act upon objects of interest ([2]). Such advances are of special import in the perspective of developing adaptable, rapidly-reconfigurable production lines. Industrial processes focusing on low-volume production of customized products have received increasing attention ([3]). Such specialized manufacturing methods require the handling of components of varying geometry and inertia. In practice today, the adjustment of production lines oftentimes remains the responsibility of specialized industrial engineers, providing expert knowledge. Such a process typically proves effort-intensive, motivating the investigation of technologies supporting its automation. A large portion of operations involved in the considered manufacturing processes require the use of a robotic arm equipped with an end-effector, supporting applications that include pick and place, tool manipulation, assembly or disassembly. These tasks commonly require that the end-effector be capable of effectively grasping objects of interest, which has motivated investigations in the area of grasp affordance learning ([4]). Results pursued are typically characterized in terms of a desirable distribution of contact points or surfaces between grasped object and robotic gripper (or equivalent representation, referred to as grasp configuration), affording a number of desirable qualities (e.g. in terms of resilience to perturbations). Work in this area has been structured along two distinct avenues of investigation, analytical and empirical approaches.
Analytical methods analyze properties of the target object, gripper, and task parameters to determine a grasp using a predefined algorithm. Commonly, such methods pursue emergence of force closure grasps, in which the grasp configuration allows the gripper to apply force on the object in all directions of motion ([5]). A grasp satisfying this criterion is intended to allow application of counteracting force against external efforts (gripper actuation system’s capabilities allowing). In [6], the authors demonstrate an approach allowing to synthesize grasps approximately satisfying closure properties for simple shapes such as polygons. The authors of [7] expand on these concepts to find all antipodal point-pairs on arbitrary 3D-objects, which can be used to synthesize grasps for prismatic grippers. This concept was expanded upon in an iterative manner, culminating in an approach capable of finding grasps containing multiple contact points, while still satisfying force closure grasp constraints ([8, 9, 10]). Recent results expand on these techniques by accounting for additional such constraints. In [11], the authors present an approach that classifies resulting grasps based on their efficacy in a cluttered environment, favoring grasps allowing to avoid collisions with nearby objects. In another extension on the approach, the result in [12] demonstrates an ability to generates grasps capable of applying mechanical stress to objects, with applications to agricultural harvests (e.g. plucking fruits from trees).
Empirical methods typically emulate previously established grasp strategies. This is commonly pursued by recording human grasps, and applying RL techniques to emulate them with the considered robotic system. The range of possible grasps are defined in an established taxonomy ([13]). In [14], the authors use a virtual reality setup to train a robot to imitate human behavior during task execution. Supervised learning has also been used to support grasp synthesis, either by observing and reproducing successful grasps ([15]), or via an annotated mix of failed and successful grasps ([16]). Recent approaches use meta-learning to pursue one-shot imitation capabilities ([17, 18]) from visual input. These models are trained using general grasp training sets (such as those in [19, 20]), then fine-tuned via demonstration. Other empirical approaches have developed grasp synthesis strategies by applying reinforcement learning techniques to physical setups featuring robotic arm, end-effector, and a series of objects of interest. The approach amounts to repeatedly attempting to grasp an object, and iteratively improving grasping behavior. Other empirical approaches rely on template matching, whereby a set of template grasps are defined for a considered set of objects. When a novel object is encountered, its shape is analyzed, similarities to known shapes identified, and an appropriate template grasp is selected. This is typically pursued by either mapping template reference frames of grasps to new objects ([21, 22]), or by simulating a considered grasp before execution ([21]).
Existing approaches suffer from a number of limitations. Quality metrics relied upon to support learning typically reflect a priori expertise, e.g. reproducing human grasps ([14, 17, 18]) or satisfying a range of criteria (e.g. force closure, [8, 9, 10]). The rationale followed is sound, as such criteria have proved to be conducive to the emergence of reliable grasping configurations. However, the approach may suffer from a lack of generalizability in situations where the considered combination of gripper and object does not conform with the a priori knowledge. More importantly, a posteriori assessment of grasp quality is typically limited to whether or not the emerging grasp is viable, with little insights into its practical resilience ([23]). The focus on grasp viability as a binary notion has additional adverse implications. In particular, to promote grasp viability, approaches oftentimes maximize end-effector joint torques or axis efforts ([24]), which is not necessarily desirable in practice, and may also color the outcome of the learning process. Further, the learning process in most instances is limited to proposing a single grasp configuration, with no information on the range of possible alternative grasps, or insights in terms of the rationale followed to reach the result ([23]), leading to a deficit of transparency and explainability (typical to RL approaches, [25]). Finally, results in the literature consider the problem from the perspective of grasp configuration. Establishing the range of such possible configurations for simple object geometries (e.g. sphere, cube) and gripper kinematics (prismatic gripper) is straightforward ([5, 6, 11, 12]). However, for more complicated object geometries and grippers, analytically establishing viable solutions rapidly becomes intractable.
To address these issues, we propose a generative approach to learning grasp affordances. We rely on the existence of a grasp synthesis policy which, given an admissible relative pose between gripper and object, defines what trajectories are applied to the gripper’s degrees of freedom (gripper base and actuated axes or joints) in pursuit of grasp emergence. Such a policy $\mathcal{G}$ provides a map from relative pose space (denoted $\mathcal{P}$ in the following) to grasp configuration space $\mathcal{C}$; that is, $\mathcal{G}:\mathcal{P}\rightarrow\mathcal{C}$. Relying on the existence of such a map, one is able to approach the problem of learning how to grasp an object by formulating grasp policies over the relative pose space $\mathcal{P}$, rather than the configuration space $\mathcal{C}$, which affords a number of benefits. Under consistent grasp policy, well-defined sets in $\mathcal{P}$ can be found to lead to consistent grasp types or behaviors. Establishing the landscape of such behavioral manifolds (i.e. sets of points $p\in\mathcal{P}$ mapping to consistent grasp types in $\mathcal{C}$ under $\mathcal{G}$, see example in Fig. 1) provides an explicit representation of the range of grasp solutions for the gripper, object, grasp policy tuple.
In the following, we explore the merit of the proposed approach. In Section II, we discuss the use of a previously introduced dynamic quality metric ([26]) to assess grasp resilience during object manipulation. Then, in Section II, we propose a simple example of grasp synthesis policy, and use it to establish the behavioral landscape in relative pose space, for a number of example objects. We further assess impact on performance of informing grasp synthesis using haptic information, and evaluate impact in terms of the size of behavioral manifolds. Section IV concludes this paper.
Figure 2: Horizontal cut of behavioral map for a cup (left), grasped with a Robotiq 3-fingered gripper, behavioral manifolds $\mathcal{A}$ through $\mathcal{D}$ map to consistent grasp behaviors (right). Contours inside manifolds show $\mu_{\rm s}$ level sets, local minima marked with yellow dots. The red line shows an example optimization trajectory from PPO (jumps in light green). Vertical cut shown bottom right, 3D perspective bottom far right. The behavioral map provides a holistic perspective of possible grasps.
II Assessing Grasp Quality
$\min\tau_{\rm s}$ $\min\tau_{\rm g}$ $\min\tau_{\rm f}$ $\min\tau_{\rm m}$
Simple Shapes
Sphere 1.48Nm 1.48Nm 1.48Nm 1.48Nm
Dumbbell 1.29Nm 1.53Nm 1.29Nm 1.31Nm
Pringles Can 1.54Nm 1.54Nm 1.55Nm 1.54Nm
Small Objects
Spoon 0.83Nm 0.84Nm 0.81Nm 0.85Nm
Fork 0.83Nm 0.84Nm 0.82Nm 0.85Nm
Screwdriver 0.93Nm 0.99Nm 1.00Nm 0.96Nm
Uneven Mass
Power Drill 1.87Nm 2.13Nm 2.31Nm 2.03Nm
Hammer 2.32Nm 2.32Nm 2.30Nm 2.33Nm
Delicate Objects
Banana 0.52Nm 0.52Nm 0.53Nm 0.52Nm
Tomato 0.51Nm 0.53Nm 0.53Nm 0.51Nm
Foam Ball 0.20Nm 0.20Nm 0.20Nm 0.20Nm
Complex Geometry
Stanford Bunny 0.88Nm 1.18Nm 1.24Nm 1.15Nm
Spiked Ball 1.46Nm 1.46Nm 1.46Nm 1.46Nm
Slippery Objects
Wet Sphere 1.94Nm 1.94Nm 1.93Nm 1.94Nm
Table 1: Minimum joint efforts required to maintain a generated grasp along a trajectory. We compare the grasps generated using the metric defined in Section II ($\tau_{\rm s}$), grasp matrix eigenvalues ($\tau_{\rm g}$), force closure grasps ($\tau_{\rm f}$), and $\tau_{\rm m}$.
In this Section, we briefly summarize the dynamic grasp metric first introduced in [27], qualitatively discuss its merit in comparison to static metrics, and explore its application to objects in the YCB data set ([28]).
II-A Dynamic Grasp Metric
[…]
$\begin{split}\ddot{\eta}_{\rm o}(t)=\begin{bmatrix}f_{\text{g}}(t)/m_{\rm o}+g\\ I^{-1}_{\text{o}}\tau_{\text{g}}(t)-2I^{-1}_{\text{o}}\Omega_{\text{o}}(t)I_{\text{o}}\omega_{\text{o}}(t)\end{bmatrix},\end{split}$ (1)
where $m_{\rm o}\in\mathds{R}$ is the object’s mass in kg, $I_{\rm o}\in\mathds{R}^{3\times 3}$ its rotational inertia matrix in kg.m, $g\in\mathds{R}^{3}$ is gravity’s acceleration in m/${\rm s}^{2}$, $f_{\text{g}}(t)$, $\tau_{\text{g}}(t)\in\mathds{R}^{3}$ describe impact of gripper-object interaction forces on the object’s dynamics (in N and N.m, respectively). Consider the following partial,
$\begin{split}\frac{\partial\ddot{\eta}_{\rm o}}{\partial\ddot{\eta}_{\rm r}}=\sum_{i=0}^{n}\frac{\partial\ddot{\eta}_{\text{o}}}{\partial f_{\text{g}}}\frac{\partial f_{\text{g}}}{\partial f_{\text{g}i}}\frac{\partial f_{\text{g}i}}{\partial\partial\ddot{\eta}_{i}}\frac{\partial\ddot{\eta}_{i}}{\partial\ddot{\eta}_{\rm r}},\end{split}$ (2)
where $f_{\text{g}i}\in\mathds{R}^{3}$ is the contact force for the $i^{\rm th}$ contact point, $i=1$, $\ldots$, $n$, and $\ddot{\eta}_{i}(t)\in\mathds{R}^{6}$ is the pose acceleration of the $i^{\rm th}$ gripper link (see [27, 26] for details). Evaluating the above partial at time instant $t_{\text{x}}$, we have
[…]
where $G(t_{\text{x}})\in\mathds{R}^{6\times 6}$ characterizes the impact of the gripper’s movements on those of the object. Minimizing relative movements between gripper and object (i.e. favoring stable relative pose) amounts to minimizing $S(t)=G(t)-I_{6}$, where $I_{6}$ is the 6-dimensional identity. Matrix $S(t)$ characterizes the extent to which the considered grasp mitigates the impact of the gripper’s movements onto the relative motion of the grasped object (see discussion of robustness to perturbations in [30]). Resilience of gripper movements along independent axes of motion can be assessed by singular value decomposition of $S(t)$, $\sigma_{S}(t)\in\mathds{R}^{6}$. Specifically, we measure grasp quality using
[…]
where $t_{0}$, $t_{\rm t}\in\mathds{R}^{+}$, are the task’s start and end times. Greater values of $\mu_{\rm s}$ correspond to better quality grasps, in terms of stability of the grasped object relative to the gripper’s base.
II-B Dynamic Metric for Grasp Learning
Static approaches ([29]) assess quality in light of a priori, fixed criteria, but do not account for the resulting behavior of the gripper-object pair during task performance. Approaches relying on such quality metrics for grasp learning may in some instances achieve undesirable results, see for illustration the example shown in Fig. 3. Vignettes in the upper-left show learned grasps for a rectangular plate by a three-digit gripper, the plot represents grasp quality (static and dynamic) as a function of maximum joint efforts. Grasp learning was conducted using Proximal Policy Optimization (PPO, [31]). The learning process is repeated, enforcing different maximum joint efforts. The grasp learned using the static metric remains consistent for all allowed joint efforts (see upper row, grasps a, b). Instead, the grasp learned using metric $\mu_{\rm s}$ changes as a function of allowed efforts. For larger efforts (to the right of the effort threshold shown in orange), the grasp learned is identical to that obtained used the static metric (grasp d). Below this threshold, i.e. when lower efforts are available, the process learns to better balance the plate, gripping it by the middle of its long side, using the double-digit side of the gripper below the plate (grasp c). In the following, we explore how the proposed dynamic metric allows to learn effective grasps, for objects from the YCB dataset, requiring smaller joint efforts.
II-C Application to YCB Dataset
Note that, as illustrated by the example in 3, increasing joint efforts commonly lead to improving value of grasp quality metrics. Accordingly, when such metrics are used to support grasp learning, they do not necessarily emphasize exploration of alternate, more advantageous contact-point locations (or grasp configurations), but may emphasize the use of increasing joint efforts. This may lead to the unintended consequence of requiring unnecessary efforts, and failing to recognize desirable grasp configurations (providing good grasp resilience, without requiring excess efforts). We investigate this notion by using a range of different grasp metrics to learn grasp affordances, for the same gripper as that shown in Fig. 1, applied to a range of objects in the YCB dataset. Then, we investigate resilience of learned grasps by reducing maximal allowed joint torques until we reach grasp failure (i.e. release of grasped object when traveling along a reference trajectory). Results are shown in Table 1 for three static grasp metrics $\mu_{\rm g}$, $\mu_{\rm f}$, and $\mu_{\rm m}$, along with the dynamic metric introduced in the previous Subsection, $\mu_{\rm s}$. The metric $\mu_{\rm g}$ maximizes grasp matrix eigenvalues as described in [32], $\mu_{\rm f}$ analyzes the object’s shape and finds the grasp that maximizes the encompassing grasp volume spanning all contact points ([29]), $\mu_{\rm m}$ uses the method described in [33] to find the grasp maximizing external perturbation resistance. For simpler geometries (sphere, cylindrical can) for which there is no special grasp affordance to learn, results are even across metrics. However, in situations where special grasp configurations may be found to promote stability, the proposed dynamic metric results in grasps resilient to lower joint torques. This is for instance the case for the hand drill in Fig. 1, where the grasp configuration around the handle, leaning against the upper part of the body, is only learned using the proposed dynamic metric, and proves more resilient than solutions reached using alternate metrics.
III Behavioral Map
The grasp configuration space $\mathcal{C}$ describes possible configurations of the gripper and object when contact occurs, and ideally, a viable grasp is formed. Definition of this space is non-unique; in some instances authors have used SE(3) ([34, 35, 36]), though it proves only sufficient to describe the gripper’s pose relative to the object. In situations in which gripper axes configuration may be unequivocally inferred from relative pose $\eta_{\rm r}$ and object geometry (e.g. as for a prismatic gripper making contact with a sphere), it may prove sufficient to characterize grasp configurations. In a general setting however, expanding the space to include gripper axes or joints coordinates $\theta_{\rm g}$ may prove necessary, leading to the following possible definition of the configuration space,
$\mathcal{C}=\left\{\;\eta_{\text{r}},\theta_{\text{g}}\;|\;n_{\text{c}}>0\right\},$ (4)
where $n_{\text{c}}$ is the number of contact points between gripper and object. Working in terms of such a set is however nontrivial. In particular, dimension of $\mathcal{C}$ may be large, depending on number of actuated degrees of freedom of the gripper, the contact constraint is not necessarily easy to work with (or even describable in a closed form for non-trivial shapes), parsing and visualizing viable regions of $\mathcal{C}$ (i.e. points in $\mathcal{C}$ that correspond to viable grasps), representing regions of $\mathcal{C}$ corresponding to better or worse grasp quality metrics, can prove less than intuitive (see one of the few attempts in [37]).
III-A Behavioral Manifolds in Pose Space
We instead propose to approach the grasp affordance learning problem from the perspective of the relative pose space
[…]
where $d_{\text{go}}\in\mathds{R}^{+}$ is the distance between gripper base and object surface, and $d_{\text{p}}\in\mathds{R}^{+}$ is a threshold distance defining the exploration radius around the object. We in addition consider a grasp synthesis policy $\mathcal{G}$, characterizing trajectories prescribed to the gripper base and actuated DoFs, intended to promote emergence of viable grasps, which operates as a mapping from relative pose to configuration space,
$\mathcal{G}:\mathcal{P}\rightarrow\mathcal{C}.$ (6)
With minimal requirements on $\mathcal{G}$, $\mathcal{P}$ is mapped to a subset of $\mathcal{C}$ under $\mathcal{G}$. Different such grasp synthesis policies are discussed in latter Subsections.
In practice, one is able to explore the relative pose space $\mathcal{P}$ and investigate outcome for different points and regions of $\mathcal{P}$ under mapping $\mathcal{G}$. We disregard points in $C$ corresponding to non-viable grasps (i.e. contact is made between part(s) of the gripper and part(s) of the object, but no effective grasp is formed), and instead focus on regions of $\mathcal{P}$ leading to the formation of viable grasps under $\mathcal{G}$. Under straightforward grasp synthesis policies, we observe the emergence of well-defined, compact (closed, bounded) sets in $\mathcal{P}$ leading to consistent grasp types. Outcome of such an exploration is represented in Fig. 2. Specifically, the diagram on the left of Fig. 2 shows a sample surface of $\mathcal{P}$ corresponding to the horizontal plane (with coordinates representing the two-dimensional position of a point of reference on the base of the gripper, gripper attitude constrained such that the gripper base is parallel to the nearest tangent of the mug body, dual-finger side of the gripper on the upper side). The object considered is a mug, the gripper is that in Fig. 1, the grasp synthesis policy that in Section III.C. Lightly dashed areas correspond to regions of space not leading to viable grasps, sets labeled $\mathcal{A}$ through $\mathcal{D}$ map to different grasp types; force closure around the cup body from $\mathcal{A}$, pinching grasp around the cup body from $\mathcal{B}$, force closure around the cup handle from $\mathcal{C}$, pinching grasp on the handle from $\mathcal{D}$ (grasp types shown in Fig. 2, top right). Emerging grasp behaviors, from each labeled set, is consistent under $\mathcal{G}$. We accordingly refer to them as behavioral manifolds. Grasp quality, as assessed by the metric in Section II, marginally varies within each set, color gradients in Fig. 2 illustrate metric gradients, level sets are represented in solid lines. Local extrema are marked with yellow dots. A projection of sets in $\mathcal{C}$ that behavioral manifolds map to under $\mathcal{G}$ are also represented in Fig. 2 (showings the position of a reference point of the gripper surface in the horizontal surface represented). Set $\mathcal{A}$ maps to the 1-dimensional set along the surface of the cup represented with a thick solid blue line, $\mathcal{B}$ maps to the light blue line on the surface of the handle, $\mathcal{C}$ and $\mathcal{D}$ to the small 2 dimensional sets located to their left, represented with thinner outlines.
The behavioral map in Fig. 2 provides a number of interesting features. In particular, it allows to visualize the range of possible grasp outcomes (both grasp type and quality, for a given $\mathcal{G}$ and metric) in a concise, self-explanatory representation. Further, consideration of behavioral manifolds provides information helpful to planning. Specifically, achieving a desired grasp type only requires reaching the relevant manifold and following the corresponding grasp synthesis. In addition, representing optimization trajectories onto the behavioral map allows to visualize and apprehend the rationale followed by the process. In Fig. 1, the trajectory generated by PPO is represented it red. In initializes in the top left corner, explores a non-viable area before entering $\mathcal{A}$. It then descends different gradients, localizes a number of local minima, follows the outline of different manifolds, jumps to different regions to escape local minima before eventually converging to a (local) minimum. The map provides explainability and transparency to the RL process.
Figure 3: Static ($\mu_{\rm g}$) and dynamic ($\mu_{\rm s}$) grasp metrics as a function of maximum joint torques, grasping a rectangular plate. Grasps learned using the static metric in the top row (grasps a, b), dynamic metric in the bottom row (c, d).
III-B Baseline Grasp Synthesis Policy
Valid grasp configurations require emergence of contact between designated parts of the gripper and object surface. In the following we consider a simple grasp policy for the three digit gripper shown in Fig. 1. From an initial relative pose at a short distance from the object, gripper base parallel to the nearest surface of the object, digits opened to their widest, such contacts are pursued by approaching the gripper base and digits towards the object’s surface. In this perspective, we define a baseline grasp synthesis policy $\mathcal{G}_{\text{c}}$. The policy relies on knowledge of the pose $\eta_{\text{r}}(t_{\text{p}})$ of the gripper’s CoM relative to the object’s CoM at initial time instant $t_{\text{p}}$. The policy then assesses (from the object’s geometry) the point on the object’s surface $x_{\text{c}}$ that is closest to the gripper base’s surface. Then, a constant force of magnitude $f_{\text{c}}$ and direction defined as that from gripper CoM to $x_{\text{c}}$ is applied to the gripper. As contact is made between gripper and object, the force is removed, and a constant closing torque $\tau_{\text{c}}$ is applied to gripper joints, until the time instant $t_{\text{c}}$ at which gripper movements have halted (under resistance from the object). The resulting gripper base pose, relative to that of the object, appended with the gripper’s joint angles, $\left(\eta_{\text{r}}(t_{\text{c}}),\theta_{\text{g}}(t_{\text{c}})\right)$ describe the grasp configuration. The proposed policy thus describes the following map,
$\mathcal{G}_{\rm{c}}:\quad\quad\mathcal{P}$$\rightarrow$ $\mathcal{C},$
$\eta_{\rm r}(t_{\rm p})$$\rightarrow$ $\eta_{\rm r}(t_{\rm c}),\;\theta_{\rm g}(t_{\rm c}).$ (7)
In the following, we adjust the previously discussed policy to exploiting proximity information made available from a haptic sensor such as that described in [38].
III-C Haptic Grasp Synthesis Policy
Figure 4: Proximity perception is used to conform shape of the gripper to that of the object, links in closer proximity move away, those further away move closer (left). Exploiting proximity information expands behavioral manifolds (right).
1 Function GraspSynthesis($\eta_{\rm r}$, $\theta_{\rm g}$, $f_{\rm s}(\cdot)$)
       /* Adjust digits so that distance to object is within tolerance */
2       while $\max\hat{d}_{{\rm s}i}>\hat{d}_{\rm a}$ do
            /* Jacobian (11) */
3             $\hat{J}_{\rm s}$ = est_finger_jacobian($\hat{d}_{{\rm s}}$, $\eta_{\rm r}$, $\theta_{\rm g}$)
4             $\theta_{\rm g}\leftarrow\hat{J}_{\rm s}^{+}\theta_{\rm d}+\theta_{\rm g}$
             /* Update digit kinematic chains */
5             foreach $i\in n$ do
6                  $\eta_{i}=k_{i}(\eta_{\rm r},\theta_{\rm g})$
      /* Reduce distance to object in a homogeneous manner for grasp emergence */
7       while $\max\hat{d}_{{\rm s}i}>\hat{d}_{\rm c}$ do
            /* Jacobian from III-C2) */
8             $J_{\rm s}$ = est_gripper_jacobian($\hat{d}_{{\rm s}}$, $\eta_{\rm r}$, $\theta_{\rm g}$)
9             $\begin{bmatrix}\eta_{\rm r}\\ \theta_{\rm g}\end{bmatrix}\leftarrow\hat{J}_{\rm s}^{+}\begin{bmatrix}\eta_{\rm d}\\ \theta_{\rm d}\end{bmatrix}+\begin{bmatrix}\eta_{\rm r}\\ \theta_{\rm g}\end{bmatrix}$
             /* Update gripper kinematic chain */
10             foreach $i\in b,[0,n)$ do
11                  $\eta_{i}=k_{i}(\eta_{\rm r},\theta_{\rm g})$
/* Return the grasp configuration */
Result: $\begin{bmatrix}\eta_{\rm r}\\ \theta_{\rm g}\end{bmatrix}$
12
Alg. 1 Shape-informed grasp synthesis policy, implementing the steps outlined in III-C1 and III-C2.
The previous grasp policy relied on knowledge of the relative pose of the object’s CoM. Hereafter, we instead adjust movements of the gripper base and digits as they approach the object, in a manner that they conform with its perceived local shape. The approach is divided in two steps. Step 1 aims to adjust gripper digits in such a manner that the distance between the surface of individual gripper links and the object’s surface nearest to the considered link are homogeneous within a given tolerance $\Delta d_{\rm s}\in\mathds{R}^{+}$. Step 2 then adjusts base pose and joint angles to reduce measured distances until a closure grasp is achieved.
III-C1 Digit Alignment
The initial condition assumes all gripper digits are in an open configuration. We evaluate the distance $d_{\rm b}(t)$ between individual links and nearest object surfaces,
$d_{{\rm s}i}(t)=f_{\rm s}(\eta_{i}(t),\eta_{\rm o}(t)),$ (8)
where $\eta_{i}(t)\in\mathds{R}^{3}$ is the pose of link $i$’s CoM, and $f_{\rm s}(\cdot)$ represents a nonlinear function computing the distance. In numerical simulation, $f_{\rm s}(\cdot)$ is provided by typical geometrical libraries, using the respective poses and 3D meshes ([39]). Specifically, we used Bullet’s integrated functions. We then define the offset distance
$\hat{d}_{{\rm s}i}(t)=d_{{\rm s}i}(t)-d_{\rm b}(t),$ (9)
where $d_{\rm b}(t)\in\mathds{R^{+}}$ is the minimum distance between base and considered object. We compute the link poses,
$\eta_{i}(t)=k_{i}(\eta_{\rm r}(t),\theta(t)),$ (10)
where $k_{i}(\cdot)$ is the links’ kinematics. Substituting (10) into (9) provides a direct relationship between joint angles and distances. We then consider the Jacobian
$\hat{J}_{\rm s}(t)=\frac{\partial\hat{d}_{\rm s}(t)}{\partial\theta_{\rm g}(t)}.$ (11)
Joint angles are adjusted in the direction of $\hat{J}_{\rm s}(t)$ until $\hat{d}_{{\rm s}n}(t)$ and $d_{\rm b}(t_{0})$ are within tolerance $\Delta d_{\rm s}$.If no joint configuration is found fulfilling the requirement, the initial pose is discarded as non-viable ($\mu_{\rm s}=0$).
III-C2 Grasp Closure
Upon completion of step 1, shape adaptation is used to pursue emergence of an enclosure grasp. We compute the Jacobian $J_{\rm s}(t)$ as the partial derivative of the distance vector $\begin{bmatrix}d_{\rm b}(t),d_{0}(t),\ldots d_{n}(t)\end{bmatrix}^{\rm T}$ with respect to base pose and joint angles, which we then adjust in the direction defined by the Jacobian. The process is repeated until a grasp emerges.
III-D Grasp Policies Comparison
Following the baseline grasps synthesis policy, the gripper’s digits simply close onto the object after the base makes contact. In situations in which the object’s geometry is simple and favorable, the policy proves successful, as shown in Fig. 3. However, in cases where the object offers a more complex surface geometry, the approach provides mixed results. For instance, small surface features offering opportunities for improved grasps can be overlooked. Fig. 4 provides an illustrative example, focusing on the opportunity for a stable force closure grasp around a cup’s handle (behavior $\mathcal{C}$ in 2). Behavioral manifolds for the baseline grasp synthesis policy are shown in a palette from grey to yellow. The policy allows emergence of both pinching behaviors, although the manifold surfaces are meaningfully reduced in comparison to those achieved using the shape-informed policy (blue palette). More significant, the manifold for handle force closure almost vanishes when using the baseline policy. In other words, in the absence of proximity information, observing emergence of this force closure grasp is unlikely. Similar results were found across different object geometries, as shown in Fig. 5. Specifically, identical relative poses led to substantially different outcomes for either policy, with the shape-informed policy providing higher quality grasps; handle force closure on cup, secure ear grasp on Stanford bunny, handle force closure on the hand drill. In addition, the emergence of larger behavioral manifolds when using the shape-informed policy was verified over all objects in the YCB dataset. Some results are shown in Table 2. In all instances, manifolds obtained from $\mathcal{G}_{\rm s}$ contain those obtained from $\mathcal{G}_{\rm c}$. The average volume increase is of the order of 7%, objects featuring non-trivial grasp affordances, such as handles, see the greatest benefit.
Baseline
Shape-aware
Figure 5: Grasps produced by $\mathcal{G}_{\rm c}$ (left) and $\mathcal{G}_{\rm s}$ (right) from the same initial relative pose; $\mathcal{G}_{\rm s}$ results in more resilient grasps.
$v_{\rm c}$ $v_{\rm s}$ $\mu_{\rm c}$ $\mu_{\rm s}$
Uneven Surface
Power Drill 136.6 158.3 0.53 0.87
Hammer 138.2 163.2 0.74 0.76
Stanford Bunny 93.5 115.5 0.64 0.69
Spiked Ball 113.9 125.4 0.43 0.43
Simple Shapes
Sphere 125.7 125.7 0.88 0.88
Cube 119.2 120.3 0.74 0.74
Dumbbell 138.2 139.2 0.94 0.94
Pringles Can 136.2 136.5 0.91 0.91
Small Objects
Spoon 93.3 93.3 0.93 0.93
Fork 93.1 93.2 0.93 0.93
Screwdriver 92.9 93.0 0.98 0.98
Table 2: Comparison of manifold volumes $v_{\rm c}$ and $v_{\rm s}$ in $[{\rm cm}^{3}]$, obtained from $\mathcal{G}_{\rm c}$ and $\mathcal{G}_{\rm s}$, respectively. Metrics $\mu_{\rm c}$ and $\mu_{\rm s}$ are provided for the best obtained grasps.
IV Conclusion
In this paper, we proposed a novel approach to grasp affordance learning, approaching the problem from the relative pose space. The method allows visualisation of possible grasp outcomes (both grasp types and quality) for different grasp synthesis policies. Different approaches to grasp synthesis are discussed, emphasizing the use of shape adaptation, supported by a haptic modality. Current efforts are investigating efficient numerical implementation of the approach. Preliminary results show convergence times 4 to 6 times faster (depending on grasped object complexity) than those observed using PPO on identical hardware. In addition, the approach is being extended to address assembly and disassembly problems, accounting from transmission of efforts, through the grasped object, onto an assembly.